OpenStudy (anonymous):

I need help with an algebra problem: The perimeter of a rectangle is 158 inches. the length exceeds the width by 57 inches. Find the length and the width.

OpenStudy (anonymous):

length = width + 57 perimeter = 2*length + 2*width substitute length. p = 2*(width+57) + 2*width = 158. solve for width. substitute width back into first equation.

OpenStudy (anonymous):

So, what is the legth & width?

OpenStudy (anonymous):

p = 2w + 114 + 2w. p = 158. 158 = 4w + 114. 158 - 114 = 4w. 44 / 4 = w = 11. l = w + 57. solve that last one yourself

OpenStudy (anonymous):

Thank you very much :)